# Mordell–Weil groups over large algebraic extensions of fields of characteristic zero

### v1

@@ 5. Kummer-faithfulness for some large algebraic extensions
Since $a$ is not a root of unity, we can take a prime number $\ell_{0}$ such that $a$ is not an $\ell$-th power in $K$ for all prime numbers $\ell\geq\ell_{0}$. Let $k$ be the algebraic closure of $\mathbb{Q}$ in $K$. By Lemma 5.7, there exists a positive integer $m_{0}$ such that $k$ and $\mathbb{Q}(\zeta_{\ell})$ for all prime numbers $\ell\geq m_{0}$ are linearly disjoint over $\mathbb{Q}$. We show that the set of prime numbers $\geq\max\{\ell_{0},m_{0},3\}$ has the desired property. The first condition is automatically satisfied since all but finitely many prime numbers belong to this set. Let $n=\ell_{1}\cdots\ell_{r}$, where $\ell_{1},\ldots,\ell_{r}$ are distinct prime numbers $\geq\max\{\ell_{0},m_{0},3\}$. Since $\mathbb{Q}(\zeta_{\ell_{1}}),\ldots,\mathbb{Q}(\zeta_{\ell_{r}})$, and $k$ are linearly disjoint over $\mathbb{Q}$, we have $[k(\zeta_{n}):k]=[\mathbb{Q}(\zeta_{n}):\mathbb{Q}]=\varphi(n)$. We also find $[K(\zeta_{n}):K]=[k(\zeta_{n}):k]=\varphi(n)$ since $K\cap k(\zeta_{n})=k$. Note that $n$ is odd as $\ell_{i}\geq 3$ for all $i$. The lemma now follows from [Lan02, Chapter VI, Theorem 9.4]. ∎
By a similar argument as in the proof of Theorem 5.2, it suffices to prove that, if we set
$S_{a}=\{\sigma\in G_{K}\mid a\in\mathbb{G}_{\mathrm{m}}(M)_{\mathrm{div}}\text{ for some finite extension }M\text{ of }\overline{K}(\sigma)\}$
for any $a\in\overline{K}^{\times}\smallsetminus\{1\}$, then we have $\mu(S_{a})=0$. If $a$ is a root of unity, then some finite extension $M$ of $\overline{K}(\sigma)$ has all roots of unity for each $\sigma\in S_{a}$. Theorem 3.1 (2-i) tells us that the set of such $\sigma\in G_{K}$ has measure zero and we obtain $\mu(S_{a})=0$.
Suppose that $a$ is not a root of unity. We continue to use some notation in the proof of Theorem 5.2. For $n\geq 1$, put
$T_{a}^{(n)}=\{\sigma\in G_{K}\mid\text{some $n$-th root of $a$ belongs to }\overline{K}(\sigma)\}$
and $\alpha^{(n)}_{1},\ldots,\alpha^{(n)}_{n}$ be all solutions of $X^{n}=a$ in $\overline{K}$. Let $T_{a}^{(n,i)}$ be the set of $\sigma\in G_{K}$ such that $\alpha^{(n)}_{i}\in\overline{K}(\sigma)$. Then we have $T_{a}^{(n)}=\bigcup_{i=1}^{n}T_{a}^{(n,i)}$ and $\bigcap_{i=1}^{n}T_{a}^{(n,i)}=G_{K(\sqrt[n]{a})}$, where $K(\sqrt[n]{a})=K(\{\alpha_{i}^{(n)}\mid 1\leq i\leq n\})$. If we prove that, for almost all $\sigma\in G_{K}$, there are infinitely many prime numbers $\ell$ such that $\sigma\notin T_{a}^{(\ell)}$, then we can verify that $a\notin\mathbb{G}_{\mathrm{m}}(M)_{\mathrm{div}}$ for any finite extension $M$ of $\overline{K}(\sigma)$ for such $\sigma$ and the proof is completed. We see that $T_{a}^{(n)}\subset G_{K(a)}$. Replacing $K$ with $K(a)$, we may assume that $a\in K^{\times}$. Let $\Lambda_{a}$ be the set of prime numbers obtained from Lemma 5.8.
@@
Remark that $K$ contains none of $\alpha_{i}^{(l)}$ for all $i$. Then
@@
[…]
@@
It follows that $\{G_{K}\smallsetminus T_{a}^{(\ell)}\}_{\ell\in\Lambda_{a}}$ is $\mu$-independent since $\{T_{a}^{(\ell)}\}_{\ell\in\Lambda_{a}}$ is so from Lemma 5.10. Using Lemma 5.9, we have
$\sum_{\ell\in\Lambda_{a}}\mu\left(G_{K}\smallsetminus T_{a}^{(\ell)}\right)=\sum_{\ell\in\Lambda_{a}}\frac{1}{\ell}=\infty.$
Applying the second Borel–Cantelli lemma (Lemma 3.4 (2)) to $\{G_{K}\smallsetminus T_{a}^{(\ell)}\}_{\ell\in\Lambda_{a}}$, we conclude that there are infinitely many $\ell\in\Lambda_{a}$ such that $\sigma\notin T_{a}^{(\ell)}$ for almost all $\sigma\in G_{K}$. ∎
@@
Let $K$, $a$, and $\Lambda_{a}$ be as in the discussion before Lemma 5.9. Then we have $\sum_{\ell\in\Lambda_{a}}\mu(T_{a}^{(\ell)})=\infty$. The first Borel–Cantelli lemma (Lemma 3.4 (1)) shows that, for almost all $\sigma\in G_{K}$, there exist infinitely many prime numbers $\ell\in\Lambda_{a}$ such that $\sigma\in T_{a}^{(\ell)}$. Lemma 2.5 implies that the torsion-free $\mathbb{Z}$-module $\mathbb{G}_{\mathrm{m}}(\overline{K}(\sigma))/\mathbb{G}_{\mathrm{m}}(\overline{K}(\sigma))_{\mathrm{tor}}$ is not free for almost all $\sigma\in G_{K}$.

### v2

@@ 5. Kummer-faithfulness for some large algebraic extensions
Since $a$ is not a root of unity, we can take a prime number $\ell_{0}$ such that $a$ is not an $\ell$-th power in $K$ for all prime numbers $\ell\geq\ell_{0}$. Let $k$ be the algebraic closure of $\mathbb{Q}$ in $K$. By Lemma 5.6, there exists a positive integer $m_{0}$ such that $k$ and $\mathbb{Q}(\zeta_{\ell})$ for all prime numbers $\ell\geq m_{0}$ are linearly disjoint over $\mathbb{Q}$. We show that the set of prime numbers $\geq\max\{\ell_{0},m_{0},3\}$ has the desired property. The first condition is automatically satisfied since all but finitely many prime numbers belong to this set. Let $n=\ell_{1}\cdots\ell_{r}$, where $\ell_{1},\ldots,\ell_{r}$ are distinct prime numbers $\geq\max\{\ell_{0},m_{0},3\}$. Since $\mathbb{Q}(\zeta_{\ell_{1}}),\ldots,\mathbb{Q}(\zeta_{\ell_{r}})$, and $k$ are linearly disjoint over $\mathbb{Q}$, we have $k(\zeta_{n}):k]=[\mathbb{Q}(\zeta_{n}):\mathbb{Q}]=\varphi(n)$. We also find $K(\zeta_{n}):K]=[k(\zeta_{n}):k]=\varphi(n)$ since $K\cap k(\zeta_{n})=k$. Note that $n$ is odd as $\ell_{i}\geq 3$ for all $i$. The lemma now follows from [Lan02, Chapter VI, Theorem 9.4]. ∎
Let $S$ be the set of $\sigma\in G_{K}$ for which the statement in Theorem 5.3 does not hold. Then what we need to show is that $S$ has measure zero in $G_{K}$. For any $a\in\overline{K}^{\times}\smallsetminus\{1\}$, let
$S_{a}=\{\sigma\in G_{K}\mid a\in\mathbb{G}_{\mathrm{m}}(\overline{K}(\sigma))_{\mathrm{div}}\}.$
Then we have
$S=\bigcup_{a\in\overline{K}^{\times}\smallsetminus\{1\}}S_{a}.$
Since the set $\overline{K}^{\times}\smallsetminus\{1\}$ is countable, it suffices to show that $S_{a}$ has measure zero for any $a\in\overline{K}^{\times}\smallsetminus\{1\}$.
Suppose that $a$ is a root of unity. Remark that a root $a\neq 1$ of unity belonging to $\mathbb{G}_{\mathrm{m}}(M)_{\mathrm{div}}$ for a field $M$ is equivalent to that $M$ contains all $\ell$-power roots of unity for some prime number $\ell$. We prove a stronger result than we need here.
Proposition 5.8.
Suppose that $K$ is a finitely generated field over $\mathbb{Q}$. Then, for almost all $\sigma\in G_{K}$, any finite extension $M$ of $\overline{K}(\sigma)$, and any prime number $\ell$, the multiplicative group $M^{\times}$ does not contain all $\ell$-power roots of unity.
Proof.
By Steps 1 and 2 in [Jar75, Section 8], we only consider the case where $K=\mathbb{Q}$. In this case, it is sufficient to show that, for any prime number $\ell$ and any positive integer $N$, the set
$E(N,\ell)=\{\sigma\in G_{\mathbb{Q}}\mid[\overline{\mathbb{Q}}(\sigma)(\zeta_{\ell^{\infty}}):\overline{\mathbb{Q}}(\sigma)]\leq N\}$
has measure zero in $G_{\mathbb{Q}}$. Here $\overline{\mathbb{Q}}(\sigma)(\zeta_{\ell^{\infty}})=\bigcup_{m=1}^{\infty}\overline{\mathbb{Q}}(\sigma)(\zeta_{\ell^{m}})$. We will prove this by showing that the measure of the set
$E(N,\ell,m)=\{\sigma\in G_{\mathbb{Q}}\mid[\overline{\mathbb{Q}}(\sigma)(\zeta_{\ell^{m}}):\overline{\mathbb{Q}}(\sigma)]\leq N\}$
tends to zero as $m\to\infty$. Suppose that $\ell$ is odd. By [Jar75, Lemma 3.1], we have
$\mu(E(N,\ell,m))=\frac{1}{\varphi(\ell^{m})}\sum_{d\leq N;d\mid\varphi(\ell^{m})}\varphi(d)\leq\frac{1}{\ell^{m-1}(\ell-1)}\sum_{d\leq N}\varphi(d).$
Since the rightmost hand side goes to zero as $m\to\infty$, we have $\mu(E(N,\ell))=0$. Suppose that $\ell=2$. We use [Jar75, Lemma 5.1] with $e=1,\theta=1/3$. Then we have
$\mu(E(N,2,m))\leq\mu(E(2^{(m-1)/3},2,m))\leq\frac{c}{2^{m/3}},$
where the first inequality holds for all $m\geq 3\log N/\log 2+1$ and $c$ is a positive constant independent of $m$. Taking $m\to\infty$ leads us to conclude that $\mu(E(N,2))=0$. ∎
Suppose that $a$ is not a root of unity. For a prime number $\ell$, put
$T_{a}^{(\ell)}=\{\sigma\in G_{K}\mid\text{some $\ell$-th root of $a$ belongs to }\overline{K}(\sigma)\}.$
We see that
$S_{a}\subset\bigcap_{\ell\text{: prime}}T_{a}^{(\ell)}.$
If we prove that, for almost all $\sigma\in G_{K}$, there exists a prime number $\ell$ such that $\sigma\notin T_{a}^{(\ell)}$, then it holds that $S_{a}$ has measure zero and the proof is completed. It is easily seen that each $\sigma\in T_{a}^{(\ell)}$ fixes $a$, that is, $T_{a}^{(\ell)}\subset G_{K(a)}$. Replacing $K$ with $K(a)$, we may assume that $a\in K^{\times}$.
Let $\alpha^{(\ell)}_{1},\ldots,\alpha^{(\ell)}_{\ell}$ be all solutions of $X^{\ell}=a$ in $\overline{K}$ and $T_{a}^{(\ell,i)}(1\leq i\leq\ell)$ be the set of $\sigma\in G_{K}$ such that $\alpha^{(\ell)}_{i}\in\overline{K}(\sigma)$. Then we have $T_{a}^{(\ell)}=\bigcup_{i=1}^{\ell}T_{a}^{(\ell,i)}$ and $\bigcap_{i=1}^{\ell}T_{a}^{(\ell,i)}=G_{K(\sqrt[\ell]{a})}$, where $K(\sqrt[\ell]{a})=K(\{\alpha_{i}^{(\ell)}\mid 1\leq i\leq\ell\})$.
Let $\Lambda_{a}$ be the set of prime numbers obtained from Lemma 5.7.
@@
Note that $K$ contains none of $\alpha_{i}^{(\ell)}$ for all $i$. Then
@@
$T_{a}^{(\ell,1)}\smallsetminus G_{K(\sqrt[\ell]{a})},\ldots,T_{a}^{(\ell,\ell)}\smallsetminus G_{K(\sqrt[\ell]{a})},\text{ and }G_{K(\sqrt[\ell]{a})}$
@@
By Lemma 5.10, the family $\{T_{a}^{(\ell)}\}_{\ell\in\Lambda_{a}}$ is $\mu$-independent. Using Lemma 5.9, we have
$\mu\left(\bigcap_{\ell\text{: prime}}T_{a}^{(\ell)}\right)\leq\mu\left(\bigcap_{\ell\in\Lambda_{a}}T_{a}^{(\ell)}\right)=\prod_{\ell\in\Lambda_{a}}\mu\left(T_{a}^{(\ell)}\right)=\prod_{\ell\in\Lambda_{a}}\left(1-\frac{1}{\ell}\right)=0.$
Therefore we conclude that there exists $\ell\in\Lambda_{a}$ such that $\sigma\notin T_{a}^{(\ell)}$ for almost all $\sigma\in G_{K}$. ∎
@@
Let $K$, $a$, and $\Lambda_{a}$ be as in the discussion before Lemma 5.9. Then we have $\sum_{\ell\in\Lambda_{a}}\mu(T_{a}^{(\ell)})=\infty$. The Borel–Cantelli lemma [FriJ23, Lemma 21.3.5] in probability theory shows that, for almost all $\sigma\in G_{K}$, there exist infinitely many prime numbers $\ell\in\Lambda_{a}$ such that $\sigma\in T_{a}^{(\ell)}$. Lemma 2.5 implies that the torsion-free $\mathbb{Z}$-module $\mathbb{G}_{\mathrm{m}}(\overline{K}(\sigma))/\mathbb{G}_{\mathrm{m}}(\overline{K}(\sigma))_{\mathrm{tor}}$ is not free for almost all $\sigma\in G_{K}$.